OpenStudy (anonymous):

If f(x) = 2x2 − 10, find f(5). 90 10 7.5 40

geerky42 (geerky42):

f(x) = f(5) so x = 5. Plug 5 in x and solve.

OpenStudy (anonymous):

You need to plug in the 5 for all the x terms in the equation If you plug it in, you get f(5)= 2(5)^2 -10 10^2 100-10 = 90 f(5)=90